The American Kennel Club recognizes 7 categories of dogs plus one for miscellaneous breeds. In 2002 Dual Championships were awarded to 141 dogs. Dual champions are dogs that show excellence in both breed standard and ability to perform the function for that breed. Out of those 141 dogs, one was awarded to a dog in the toy class and 3 to dogs in the herding breeds. The remaining Dual Championships were awarded to dogs in the sporting breeds and the hound breeds. There were 93 more Dual Championships awarded to dogs in the hound breeds than the sporting breeds. (Source: American Kennel Club)
Let s = number of sporting breeds awarded Dual Championships
Let h = number of hound breeds awarded Dual Championships.
1. Write a system of equations that could be used to determine the number of Dual Championships awarded to sporting breeds and hound breeds.

Answer by stanbon(48568)

Out of those 141 dogs, one was awarded to a dog in the toy class and 3 to dogs in the herding breeds. The remaining Dual Championships were awarded to dogs in the sporting breeds and the hound breeds. There were 93 more Dual Championships awarded to dogs in the hound breeds than the sporting breeds. (Source: American Kennel Club)
Let s = number of sporting breeds awarded Dual Championships
Let h = number of hound breeds awarded Dual Championships.
1. Write a system of equations that could be used to determine the number of Dual Championships awarded to sporting breeds and hound breeds.
------------
s+h=1411-1-3=137
h=s+93
---------
These could be rewritten as:
s+h=137
s-h=-93
Cheers,
Stan H.